# Same Task, More Tokens: the Impact of Input Length on the Reasoning Performance of Large Language Models

### v1

@@ Abstract
This paper explores the impact of extending input lengths on the capabilities of Large Language Models (LLMs). Despite LLMs advancements in recent times, their performance consistency across different input lengths is not well understood. We investigate this aspect by introducing a novel QA reasoning framework, specifically designed to assess the impact of input length. We isolate the effect of input length using multiple versions of the same sample, each being extended with padding of different lengths, types and locations. Our findings show a notable degradation in LLMs’ reasoning performance at much shorter input lengths than their technical maximum. We show that the degradation trend appears in every version of our dataset, although at different intensities. Additionally, our study reveals that traditional perplexity metrics do not correlate with performance of LLMs’ in long input reasoning tasks. We analyse our results and identify failure modes that can serve as useful guides for future research, potentially informing strategies to address the limitations observed in LLMs.
@@
[…]
@@ 2 Desired Data Properties
    To avoid model reliance on parametric knowledge rather than on the text, and to avoid data contamination Jacovi et al. (2023) the question and supporting relevant spans should consist of novel facts not seen in training. [^3] Models often answer the question correctly even if none, or only one of the required supporting facts is present in its input. We discuss this further in Appendix A.
@@ 3 FLenQA
We release the dataset to support future studies of reasoning and long input performance. Details and statistics of the tasks appear in Appendix B.
@@
Throughout the text, key paragraphs are typeset in red, the supporting sentences within them in darker red, and the optional prefixes and suffixes in black. The full prompts used for each dataset are in Appendix D.
@@ 3.2 The tasks
Julie Baker is younger than Julian Barton. This is a fact that remains constant, unchanging like the northern star. It’s a truth that is as clear as day that she ...
@@
John’s living room is marble-floored, a reality that is as intrinsic to the building as its very foundations. The moment ...
@@
We employ the task formulation from Ruletaker Clark et al. (2021), a benchmark designed for theorem proving within texts that present explicit logical theories in natural language. Each instance consists of a logical rule, two sentences each introcuding a fact, and a question over the rule and facts. [^5] Initial experiments revealed that most LLMs still struggle with instances involving multiple rules or more than two facts. Our Simplified Ruletaker task consists of generated samples that fit these criteria.
@@
Erin is good. Erin was always known for how good he is. His goodness appears on all matters of life ...
@@ 3.3.1 Background Texts
To get background text that is similar to the key paragraphs, we pad using paragraphs sampled from other base instances of the same task. To avoid creating contradictions, we exclude paragraphs that contain entities appearing in the key paragraphs. This padding therefore does not produce adversarial or ambiguous versions of the samples.
@@ 3.4 Base instances are answerable
We estimate the baseline accuracy by evaluating the LLMs on the minimal text of each sample in the dataset that includes only the question and the key paragraphs relevant to it. We found that even when using non-CoT prompting, four out of the five models achieve high accuracy (>0.89). The lowest performing model (GPT3.5) achieve high enough accuracy for degradation to be observable (0.77). Full results can be found in Appendix C.
@@ 4 Main Experiments
We report average accuracies over all three tasks, and maintain the same setup (prompt, temperature, etc.) over all input lengths. We evaluate five recent capable LLMs: GPT4, GPT3.5, Gemini-Pro, Mistral 70B and Mixtral 8x7B. See Appendix E for a detailed breakdown of our setup parameters.
@@ 4.1 Impact of Length and Location
We first look into the extreme case where only relevant tokens are added (“duplicate padding”). Shi et al. (2023) Demonstrate that appending irrelevant texts to the input of a reasoning task (GSM-8K Cobbe et al. (2021)) reduces model performance substantially. We isolate the effect of relevance by testing a setting in which the padding is duplications of the exact text of the key paragraphs. In this setup, the LLMs are not required to “search" the input to find the key paragraphs, so any bias towards any position becomes irrelevant Liu et al. (2023b). Also, any difficulty that might be imposed by the distance between the key paragraphs also becomes irrelevant. Hence, we expect that there will be no degradation in performance. Surprisingly, the Results shown in Figure 3, reveal that even in this setup length does play a factor, and accuracy decreases with length for all models.
@@
Figure 4: Effect of key paragraphs positions on accuracy. Accuracy decreases as input length grows regardless of where the key paragraphs are placed within the input. Each point reflects 300 samples. Results for all models appear in Appendix E
The results in Figure 4 show a significant drop in accuracy as length increase beyond 500 tokens. For most models, adjacency of key paragraphs produces higher accuracy, and when the key paragraphs appear last, accuracy is often highest (suggesting recency bias).
@@ 5 Correlation with Next Word Prediction
Perplexity is used as the main benchmark to show that models utilize their entire input Anil et al. (2023); Jiang et al. (2024). However, it was shown that performance on downstream tasks does not necessarily correlate with model perplexity (Liu et al., 2023a; Xia et al., 2022; Tay et al., 2022). Here, we will use the flexibility of our dataset to understand the correlation between perplexity and reasoning accuracy.
@@ 6 Does Chain of Thought Help?
The results show (Figure 1) that CoT has different effects on different LLMs, and overall does not mitigate the drop in performance due to length. In most cases (GPT4, Mixtral 8x7B, Mistral 70B and GPT3.5) it improves performance, but only in GPT4 it has an increased effect as length increases, making it a limited mitigation technique. In the case of Gemini-Pro, we see that CoT decrease performance as input length is increased, even though it increase performance on short length.
The full results of the CoT prompting over all tasks and setups can be found in Appendix E.
@@ 7 Length-induced Failure modes
All of the samples in the dataset can be answered with either "True" or "False", as instructed in our prompts (Appendix D). However, some of LLMs responded with a refusal answer the question, often preceded by a sentence such as "There is not enough information in the text". This tendency grows as the input length increases, indicating a failure to comply to the instruction that specified a clear choice between "True" and "False". The trend is demonstrated in figure 7, and results over all models in Appendix 16.
@@
When using Chain-of-Thought prompting, some LLMs were much more likely to output the final true/false answer before the expected reasoning steps, as inputs grow longer. In recent work, Kojima et al. 2022 found that when models are elicited to provide the reasoning steps after the answer their performance does not increase (as expected when using auto-regressive models that only attend to earlier tokens). This can be viewed as a case of failing to follow prompt instructions (see prompt instructions in Appendix D) as the length increases. In testing, we found that incorrect responses are statistically dependent on the occurrence of answers before the reasoning steps [^9] Corresponding odds-ratio is 3.643 with $p<0.001$ obtained through Fisher exact test..
@@
We measure this by computing the coverage of the relevant text (the key sentences in each sample) in the models’ “steps" part of the outputs (details in Appendix D.4). We find that in most models, the ability to locate the relevant text within the input decreases as the input length gets longer. We found incorrect responses were statistically dependent on the incomplete coverage of the facts [^10] Corresponding odds-ratio is 3.138 with $p<0.001$ obtained through Fisher exact test..
@@ 8 Related Work
The evaluation of LLMs on long inputs has followed two distinct pathways: benchmarks of downstream tasks and next word prediction. In the realm of benchmarks, studies proposed datasets of long input samples that can be used to evaluate models Shaham et al. (2023, 2022); An et al. (2023b, a). Those datasets are curated over inputs of different, but fixed, length. This approach, while straightforward, limits our ability to inputs of varying lengths, posing a challenge in understanding the true impact of input length on model performance. On the other hand, next word prediction evaluations do offer an insights into how models handle inputs of different lengths (like done in Anil et al. 2023; Jiang et al. 2024). However, the correlation of this task with downstream performance was found not consistent (Liu et al., 2023a; Xia et al., 2022; Tay et al., 2022). In this paper we reproduce this finding with respect to extended length.
@@ 9 Discussion
Because of the nature of behavioral testing, the observed drop in performance with varying input lengths remains unexplained; because of lack of access to many of the models, we suspect this direction will continue to be limited. Secondly, our approach aimed to create a universally applicable test across different LLMs, leading to the selection of tasks that cater to the lowest common denominator. This approach potentially overlooks the nuanced performance differences in more complex reasoning tasks (e.g 5 key paragraphs), where, for instance, stronger models might exhibit performance degradation at shorter input lengths compared to what our findings suggest. Additionally, we focused on a subset of reasoning task types which may differ behaviourally from other types. Finally, our study did not test the distance between key paragraphs, leaving an aspect of LLM performance unexplored that we leave for future research.
@@ Acknowledgements
Appendix A Training Contamination in Reasoning Tasks
Data contamination is a major concern when evaluating models (Sainz et al., 2023; Jacovi et al., 2023). Ensuring that a task requires reasoning across multiple text spans is a stronger requirement then a task that requires multi hop reasoning (§2). Evaluating models on questions they answer using parametric knowledge prevents us from assessing their reasoning capabilities. Datasets originating from internet sources are especially vulnerable to contamination, thereby undermining the evaluation of a model’s capacity to reason over novel facts.
Furthermore, it was shown that small models can answer existing reasoning dataset when given one some of the parts that were claimed to be required for the task (Chen and Durrett, 2019; Min et al., 2019). This raises question on how well those dataset can be used to evaluate the effect of length, where we require that the reasoning will be done on multiple parts of the text.
We conclude that parametric knowledge should be accounted for when evaluating text-based reasoning capabilities. In this work we introduced FlenQA, which is composed of novel generated data to make sure that reasoning over the input is required.
Appendix B Datasets
@@ B.3 People in Rooms (PIR)
Appendix C Base Instances Result
Model Prompt MonoRel PIR Ruletaker*
GPT3.5 Turbo Direct 0.77 0.81 0.74
CoT 0.86 0.88 0.88
GPT4 Turbo Direct 1.00 1.00 0.98
CoT 1.00 1.00 0.97
Gemini Pro Direct 0.84 1.00 0.92
CoT 0.88 0.96 0.97
Mistral 70B Direct 0.99 1.00 0.73
CoT 1.00 1.00 0.89
Mixtral 8x7B Direct 0.92 0.97 0.80
CoT 0.86 0.97 0.93
Appendix D Full Evaluation Setup
D.1 Prompts
@@ D.2 Parameters
D.3 Locating the answer in models’ replies
@@
D.4 Evaluating the coverage of key facts in CoT
@@
Appendix E Full results

### v2

@@ Abstract
This paper explores the impact of extending input lengths on the capabilities of Large Language Models (LLMs). Despite LLMs advancements in recent times, their performance consistency across different input lengths is not well understood. We investigate this aspect by introducing a novel QA reasoning framework, specifically designed to assess the impact of input length. We isolate the effect of input length using multiple versions of the same sample, each being extended with padding of different lengths, types and locations. Our findings show a notable degradation in LLMs’ reasoning performance at much shorter input lengths than their technical maximum. We show that the degradation trend appears in every version of our dataset, although at different intensities. Additionally, our study reveals that the traditional metric of next word prediction correlates negatively with performance of LLMs’ on our reasoning dataset. We analyse our results and identify failure modes that can serve as useful guides for future research, potentially informing strategies to address the limitations observed in LLMs.
@@
Mosh Levythanks: These authors contributed equally to this work.1  Alon Jacoby11footnotemark: 11  Yoav Goldberg1,2 1Bar-Ilan University      2Allen Institute for AI {moshe0110, alonj4}@gmail.com
@@ 2 Desired Data Properties
    The question and supporting relevant spans should consist of novel facts not seen in training. Ensuring that a task requires reasoning across multiple text spans is a stronger requirement then a task that requires multi hop reasoning. It was shown that models can answer existing reasoning dataset when given one some of the parts that were claimed to be required for the task (Chen and Durrett, 2019; Min et al., 2019). To avoid model reliance on parametric knowledge when we expect a reasoning process to be done (i.e to avoid data contamination Jacovi et al. (2023); Sainz et al. (2023)), we desire that an evaluation aimed to test reasoning capabilities will require reasoning over texts that were not previously available.
@@ 3 FLenQA
We release the dataset and the code to generate it from scratch to support future studies of reasoning and long input performance. Generating the dataset from scratch can be used to prevent data contamination in future evaluation. Details and statistics of the tasks appear in Appendix A.
@@
Throughout the text, key paragraphs are typeset in red, the supporting sentences within them in darker red, and the optional prefixes and suffixes in black. The full prompts used for each dataset are in Appendix B.
@@ 3.2 The tasks
Julie Baker is younger than Julian Barton.
This is a fact that remains constant, unchanging like the northern star. It’s a truth that is as clear as day that she ...
@@
John’s living room is marble-floored
, a reality that is as intrinsic to the building as its very foundations. The moment ...
@@
We employ the task formulation from Ruletaker Clark et al. (2021), a benchmark designed for theorem proving within texts that present explicit logical theories in natural language. Each instance consists of a logical rule, two sentences each introducing a fact, and a question over the rule and facts. [^4] Initial experiments revealed that most LLMs still struggle with instances involving multiple rules or more than two facts. Our Simplified Ruletaker task consists of generated samples that fit these criteria.
@@
Erin is good.
Erin was always known for how good he is. His goodness appears on all matters of life ...
@@ 3.3.1 Background Texts
To get background text that is similar to the key paragraphs, we pad using paragraphs sampled from other base instances of the same task. To avoid creating contradictions, we exclude paragraphs that contain entities appearing in the key paragraphs. This padding therefore does not produce adversarial or ambiguous versions of the samples. This type of padding creates an input that resembles the RAG setup, where the input is composed of independent texts from a similar source (Mao et al., 2020).
@@ 3.4 Base instances are answerable
We estimate the baseline accuracy by evaluating the LLMs on the minimal text of each sample in the dataset that includes only the question and the key paragraphs relevant to it. The results of the base instances are brought in 3.4. We found that even when using non-CoT prompting, four out of the five models achieve high accuracy (>0.89). The lowest performing model (GPT3.5) achieve high enough accuracy for degradation to be observable (0.77).
Model Prompt MonoRel PIR Ruletaker*
GPT3.5 Direct 0.77 0.81 0.74
CoT 0.86 0.88 0.88
GPT4 Direct 1.00 1.00 0.98
CoT 1.00 1.00 0.97
Gemini Pro Direct 0.84 1.00 0.92
CoT 0.88 0.96 0.97
Mistral Medium Direct 0.99 1.00 0.73
CoT 1.00 1.00 0.89
Mixtral 8x7B Direct 0.92 0.97 0.80
CoT 0.86 0.97 0.93
@@ 4 Main Experiments
We report average accuracies over all three tasks, and maintain the same setup (prompt, temperature, etc.) over all input lengths. We evaluate five recent capable LLMs: GPT4, GPT3.5, Gemini-Pro, Mistral Medium and Mixtral 8x7B. We consider an output where no answer was mentioned (e.g "I don’t know") as incorrect. See Appendix C for a detailed breakdown of our setup parameters.
@@ 4.1 Impact of Length and Location
We first look into the extreme case where only relevant tokens are added (“duplicate padding”). Shi et al. (2023) Demonstrate that appending irrelevant texts to the input of a reasoning task (GSM-8K Cobbe et al. (2021)) reduces model performance substantially. We isolate the effect of relevance by testing a setting in which the padding is duplications of the exact text of the key paragraphs. In this setup, the LLMs are not required to “search" the input to find the key paragraphs, so any bias towards any position becomes irrelevant Liu et al. (2023b). Also, any difficulty that might be imposed by the distance between the key paragraphs also becomes irrelevant. Hence, we expect that there will be no degradation in performance. The Results shown in Figure 3, reveal that even in this setup length does play a factor, and accuracy decreases with length for all models. We consider these results surprising: duplicated texts are an artificial setup which is arguably the best case scenario of long inputs, as the information is constantly repeated and there is no distracting text. In more natural cases, most of the input is irrelevant to question asked. We test this setup in the next section.
@@
Figure 4: Accuracy decreases as input length grows regardless of where the key paragraphs are placed within the input. Each point reflects 300 samples. Results for all models appear in Appendix C
The results in Figure 4 show a significant drop in accuracy as length increase beyond 500 tokens. For most models, adjacency of key paragraphs produces higher accuracy, and when the key paragraphs appear last, accuracy is often highest (suggesting recency bias). We also find that some models perform worse when the key paragraphs are in the middle, similarly to what was found in the extraction task studied recently (Liu et al., 2023b).
@@ 5 Correlation with Next Word Prediction
Perplexity is used as the main benchmark to show that models utilize their entire input Anil et al. (2023); Jiang et al. (2024); Ding et al. (2024). However, it was shown that performance on downstream tasks does not necessarily correlate with model perplexity (Liu et al., 2023a; Xia et al., 2022; Tay et al., 2022). Here, we will use the flexibility of our dataset to understand the correlation between perplexity and reasoning accuracy.
@@ 6 Does Chain of Thought Help?
The results show (Figure 1) that CoT has different effects on different LLMs, and overall does not mitigate the drop in performance due to length. In most cases (GPT4, Mixtral 8x7B, Mistral Medium and GPT3.5) it improves performance, but only in GPT4 it has an increased effect as length increases, making it a limited mitigation technique. In the case of Gemini-Pro, we see that CoT decrease performance as input length is increased, even though it increase performance on short length.
The full results of the CoT prompting over all tasks and setups can be found in Appendix C.
@@ 7 Length-induced Failure modes
All of the samples in the dataset can be answered with either "True" or "False", as instructed in our prompts (Appendix B). However, some of LLMs responded with a refusal answer the question, often preceded by a sentence such as "There is not enough information in the text". This tendency grows as the input length increases, indicating a failure to comply to the instruction that specified a clear choice between "True" and "False". The trend is demonstrated in figure 7, and results over all models in Appendix 16.
@@
When using Chain-of-Thought prompting, some LLMs were much more likely to output the final true/false answer before the expected reasoning steps, as inputs grow longer. In recent work, Kojima et al. 2022 found that when models are elicited to provide the reasoning steps after the answer their performance does not increase (as expected when using auto-regressive models that only attend to earlier tokens). This can be viewed as a case of failing to follow prompt instructions (see prompt instructions in Appendix B) as the length increases. In testing, we found that incorrect responses are statistically dependent on the occurrence of answers before the reasoning steps [^8] Corresponding odds-ratio is 3.643 with $p<0.001$ obtained through Fisher exact test..
@@
We measure this by computing the coverage of the relevant text (the key sentences in each sample) in the models’ “steps" part of the outputs (details in Appendix B.4). We find that in most models, the ability to locate the relevant text within the input decreases as the input length gets longer. We found incorrect responses were statistically dependent on the incomplete coverage of the facts [^9] Corresponding odds-ratio is 3.138 with $p<0.001$ obtained through Fisher exact test..
@@ 8 Related Work
The evaluation of LLMs on long inputs has followed two distinct pathways: benchmarks of downstream tasks and next word prediction. In the realm of benchmarks, studies proposed datasets of long input samples that can be used to evaluate models Shaham et al. (2023, 2022); An et al. (2023b, a); Bai et al. (2023). Those datasets are curated over inputs of different, but fixed, length. This approach, while straightforward, limits our ability to inputs of varying lengths, posing a challenge in understanding the true impact of input length on model performance. On the other hand, next word prediction evaluations do offer an insights into how models handle inputs of different lengths (like done in Anil et al. 2023; Jiang et al. 2024). However, the correlation of this task with downstream performance was found not consistent (Liu et al., 2023a; Xia et al., 2022; Tay et al., 2022). In this paper we reproduce this finding with respect to extended length.
@@ 9 Discussion
Because of the nature of behavioral testing, the observed drop in performance with varying input lengths remains unexplained; because of lack of access to many of the models, we suspect this direction will continue to be limited. Secondly, our approach aimed to create a universally applicable test across different LLMs, leading to the selection of tasks that cater to the lowest common denominator. This approach potentially overlooks the nuanced performance differences in more complex reasoning tasks (e.g 5 key paragraphs), where, for instance, stronger models might exhibit performance degradation at shorter input lengths compared to what our findings suggest. We focused on a subset of reasoning task types which may differ behaviourally from other types. Moreover, in order to extend the key sentences to key paragraphs, we employed GPT4 which may introduced some level bias to the text that surrounded the text required to the reasoning task (that was generated without GPT4). Finally, our study did not test the distance between key paragraphs, leaving an aspect of LLM performance unexplored that we leave for future research.
@@ Acknowledgements
This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme, grant agreement No. 802774 (iEXTRACT).
The authors would like to thank Uri Katz, Royi Rassin and Natalie Shapira for illuminating discussions and comments.
Appendix A Datasets
@@ A.3 People in Rooms (PIR)
Appendix B Full Evaluation Setup
B.1 Prompts
@@ B.2 Parameters
B.3 Locating the answer in models’ replies
@@
B.4 Evaluating the coverage of key facts in CoT
@@
Appendix C Full results